# Estimates for Dirichlet Eigenvalues of the Schrödinger operator with the Kronig-Penney Model

### v1

@@ 2 On the Sharp Asymptotic Formulas
$D_{2}(n,p)=\frac{i}{4\pi n}\int_{0}^{\pi}p(x)G(x,n)e^{-i2nx}\,dx$
$\qquad=\frac{i}{4\pi n}\biggl(\int_{0}^{\pi-c}p(x)G(x,n)e^{-i2nx}\,dx+\int_{\pi-c}^{\pi}p(x)G(x,n)e^{-i2nx}\,dx\biggr).$
We first calculate the terms in $D_{1}(n,p)$. Using $(b-a)c=b\pi$ or $ac=b(c-\pi)$ by (10), $p_{2n}=q_{-2n}=\dfrac{i(b-a)(e^{i2nc}-1)}{2\pi n}$, and (37), we have
@@
$\int_{0}^{\pi-c}p(x)Q(x,n)e^{-i2nx}\,dx=\int_{0}^{\pi-c}b\biggl(\dfrac{a}{i2n}(e^{i2nx}-1)-q_{-2n}x\biggr)e^{-i2nx}\,dx$
$\qquad=\dfrac{ab}{i2n}\int_{0}^{\pi-c}(1-e^{-i2nx})\,dx-bq_{-2n}\int_{0}^{\pi-c}xe^{-i2nx}\,dx$
$\qquad=\dfrac{ab}{i2n}\biggl(\pi-c+\dfrac{e^{i2nc}-1}{i2n}\biggr)+\dfrac{bq_{-2n}}{i2n}\biggl((\pi-c)e^{i2nc}+\dfrac{e^{i2nc}-1}{i2n}\biggr),$
@@
$\int_{\pi-c}^{\pi}p(x)Q(x,n)e^{-i2nx}\,dx=\int_{\pi-c}^{\pi}a\biggl(\dfrac{b}{i2n}(e^{i2nx}-1)-q_{-2n}x+\pi q_{-2n}\biggr)e^{-i2nx}\,dx$
$\qquad=\dfrac{ab}{i2n}\int_{\pi-c}^{\pi}(1-e^{-i2nx})\,dx-aq_{-2n}\int_{\pi-c}^{\pi}xe^{-i2nx}\,dx+a\pi q_{-2n}\int_{\pi-c}^{\pi}e^{-i2nx}\,dx$
$\qquad=\dfrac{ab}{i2n}\biggl(c-\dfrac{e^{i2nc}-1}{i2n}\biggr)+\dfrac{aq_{-2n}}{i2n}\biggl(\pi-(\pi-c)e^{i2nc}-\dfrac{e^{i2nc}-1}{i2n}\biggr)+\dfrac{a\pi q_{-2n}}{i2n}(e^{i2nc}-1).$
Therefore,
$\qquad=\frac{i}{4\pi n}\biggl(\frac{ab\pi}{i2n}+\frac{(a-b)(e^{i2nc}-1)q_{-2n}}{4n^{2}}-\frac{a\pi(e^{i2nc}-1)q_{-2n}}{i2n}\biggr)-\frac{iQ_{n,0}p_{2n}}{4n}$
and substituting $p_{2n}=q_{-2n}=\dfrac{i(b-a)(e^{i2nc}-1)}{2\pi n}$ in the last expression, we have
$D_{1}(n,p)=\frac{ab}{8n^{2}}-\frac{i(b-a)(b+3a)(e^{i2nc}-1)^{2}}{32\pi n^{3}}+\frac{(b-a)^{2}(e^{i2nc}-1)^{2}}{16\pi^{2}n^{4}}.$ (41)
Now let us calculate the terms in $D_{2}(n,p)$:
$\int_{0}^{\pi-c}p(x)G(x,n)e^{-i2nx}\,dx=\int_{0}^{\pi-c}b\biggl(\dfrac{a}{i2n}(e^{i2nx}-1)+\dfrac{\pi}{2}q_{-2n}x(e^{ix}-1)\biggr)e^{-i2nx}\,dx$
$\qquad=\dfrac{ab}{i2n}\int_{0}^{\pi-c}(1-e^{-i2nx})\,dx+b\dfrac{\pi}{2}q_{-2n}\int_{0}^{\pi-c}xe^{-i(2n-1)x}\,dx-b\dfrac{\pi}{2}q_{-2n}\int_{0}^{\pi-c}xe^{-i2nx}\,dx$
$\qquad=\dfrac{ab}{i2n}\biggl(\pi-c+\dfrac{e^{i2nc}-1}{i2n}\biggr)+\dfrac{b\pi q_{-2n}}{i2(2n-1)}\biggl((\pi-c)e^{i(2n-1)c}+\dfrac{e^{i(2n-1)c}+1}{i(2n-1)}\biggr)$
$\qquad+\dfrac{b\pi q_{-2n}}{i4n}\biggl((\pi-c)e^{i2nc}+\dfrac{e^{i2nc}-1}{i2n}\biggr)$
and
$\int_{\pi-c}^{\pi}p(x)G(x,n)e^{-i2nx}\,dx=\int_{\pi-c}^{\pi}a\biggl(\dfrac{b}{i2n}(e^{i2nx}-1)+\dfrac{\pi}{2}q_{-2n}x(e^{ix}-1)+\pi q_{-2n}\biggr)e^{-i2nx}\,dx$
$\qquad=\dfrac{ab}{i2n}\int_{\pi-c}^{\pi}(1-e^{-i2nx})\,dx+a\dfrac{\pi}{2}q_{-2n}\int_{\pi-c}^{\pi}xe^{-i(2n-1)x}\,dx-a\dfrac{\pi}{2}q_{-2n}\int_{\pi-c}^{\pi}xe^{-i2nx}\,dx+a\pi q_{-2n}\int_{\pi-c}^{\pi}e^{-i2nx}\,dx$
$\qquad=\dfrac{ab}{i2n}\biggl(c-\dfrac{e^{i2nc}-1}{i2n}\biggr)+\dfrac{a\pi q_{-2n}}{i2(2n-1)}\biggl(\pi-(\pi-c)e^{i(2n-1)c}+\dfrac{-e^{i(2n-1)c}+1}{i(2n-1)}\biggr)$
$\qquad+\dfrac{a\pi q_{-2n}}{i4n}\biggl(\pi-(\pi-c)e^{i2nc}-\dfrac{e^{i2nc}-1}{i2n}\biggr)+\dfrac{a\pi q_{-2n}}{i2n}(e^{i2nc}-1)$
give
$D_{2}(n,p)=\frac{i}{4\pi n}\int_{0}^{\pi}p(x)G(x,n)e^{-i2nx}\,dx$
$\qquad=\frac{i}{4\pi n}\biggl(\frac{ab\pi}{i2n}+\frac{a\pi^{2}(1-e^{i(2n-1)c})q_{-2n}}{i2(2n-1)}-\frac{\pi(b+a+(b-a)e^{i(2n-1)c})q_{-2n}}{2(2n-1)^{2}}$
$\qquad+\frac{a\pi^{2}(1-e^{i2nc})q_{-2n}}{i4n}+\frac{\pi(b-a)(1-e^{i2nc})q_{-2n}}{8n^{2}}+\dfrac{a\pi(e^{i2nc}-1)q_{-2n}}{i2n}\biggr).$
Using $q_{-2n}=\dfrac{i(b-a)(e^{i2nc}-1)}{2\pi n}$, we obtain
@@
Thus, by (41), (42), and Lemma 1, we write
$D(n,p)=D_{1}(n,p)+D_{2}(n,p)=\frac{ab}{4n^{2}}-\frac{i(b-a)(b+3a)(e^{i2nc}-1)^{2}}{32\pi n^{3}}+\frac{(b-a)^{2}(e^{i2nc}-1)^{2}}{16\pi^{2}n^{4}}$
$\qquad+\frac{i(b-a)(1-e^{i(2n-1)c})(e^{i2nc}-1)}{16n^{2}(2n-1)}+\frac{(b^{2}-a^{2}+(b-a)^{2}e^{i(2n-1)c})(e^{i2nc}-1)}{16\pi n^{2}(2n-1)^{2}}$
$\qquad-\frac{ia(b-a)(e^{i2nc}-1)^{2}}{32n^{3}}+\frac{(b-a)^{2}(e^{i2nc}-1)^{2}}{64\pi n^{4}}+\frac{ia(b-a)(e^{i2nc}-1)^{2}}{16\pi n^{3}}.$
and simplifying a bit
@@
Now we calculate $B_{1,n}(n^{2})$ defined by (22) in Lemma 2 . Expressing $Q(x)$ defined in Lemma 2 for the Kronig-Penney potential as
$Q(x)=\begin{cases}ax&\text{if }x\in[0,c],\\[5.69054pt] bx-b\pi&\text{if }x\in(c,\pi]\end{cases},$
@@
Therefore, by
$C_{k}=\dfrac{1}{\pi}{\displaystyle\int_{0}^{\pi}}q(x)\cos kxdx=\frac{a-b}{\pi k}\sin kc,$
(40), (43), and (44), we can calculate:
$-C_{2n}-\frac{1}{\pi}\int_{0}^{\pi}Q^{2}(x)\cos 2nxdx+D(n,q)+2\operatorname{Re}D(n,p)$
$\qquad=-\frac{a-b}{2\pi n}\sin 2nc+\frac{ab\cos 2nc}{2n^{2}}+\frac{ab}{4n^{2}}+O\bigl(\frac{1}{n^{3}}\bigr).$
@@
$\lambda_{n}=n^{2}+\frac{b-a}{2\pi n}\sin 2nc+\frac{ab(2\cos 2nc+1)}{4n^{2}}+O\bigl(\frac{1}{n^{3}}\bigr).$
@@ 3 Estimates for the Small Eigenvalues
in the set $I_{n}:=[n^{2}-M,n^{2}+M]$, where $M=\max\{|a|,b\}$, $C_{k}=\dfrac{1}{\pi}{\displaystyle\int_{0}^{\pi}}q(x)\cos kxdx$ and the infinite series $a_{k}$ is defined by (5), if the following condition holds:
@@
To consider the $n$th eigenvalue $\lambda_{n}$ for $n>1,$ we assume that the condition M<(2n1)/2M<(2n-1)/2 holds. In case $n=1$, we assume that the condition $M\leq 1/2$ holds.
Then, to use numerical methods we take finite sums instead of the infinite series in the equations obtained. To approximate the roots of the equation (45), we use the fixed point iteration. It can also be used the Newton-Raphson method but in this case it is necessary to compute the derivative of the function $K_{n}(\lambda)$ defined by (55). Then, using the Banach fixed point theorem, we prove that each of these equations containing the finite sums has a unique solution in the appropriate set $I_{n}$ (see Theorem 4). Moreover, we give error analysis (see Theorem 4 and Theorem 5) and present a numerical example.
Now, we state some preliminary facts. It is well known that the spectrum of the operator $L(q)$ is discrete and for large enough $n$, there is one eigenvalue (counting with multiplicity) in the neighborhood of $n^{2}$. See the basic and detailed classical results in [1, 2, 4, 5] and references therein. The eigenvalues of the operators $L(0)$ are $n^{2}$, for $n\in\mathbb{Z}^{+}$ and all the eigenvalues of $L(0)$ are simple.
It is also known that [6]
$|\lambda_{n}-n^{2}|\leq M,$
for $n\geq 1$. Therefore, we have
@@
for $n\geq 1$. If $k\neq\pm n$, then
$|\lambda_{n}-k^{2}|\geq|n^{2}-k^{2}|-M=|n-k||n+k|-M$
$\qquad\geq(2n-1)-M,$ (47)
for $n\geq 1$ and under the assumptions about $M$ given in Condition 1. In particular, if $n=1$, we have $|\lambda_{1}|\leq 1+M$ and
$|\lambda_{1}-k^{2}|\geq||\lambda_{1}|-k^{2}|\geq 4-|\lambda_{1}|\geq 3-M,$
for $|k|\geq 2$. Besides, if $n\geq 2$, we have $|\lambda_{n}|\geq|\lambda_{2}|\geq 4-M$ and
$|\lambda_{n}-k^{2}|\geq||\lambda_{2}|-k^{2}|\geq|\lambda_{2}|-1\geq 3-M,$
for $k\neq\pm n$.
@@
We remind that, since the potential $q$ is of the form (9), we have
$C_{k}=\frac{a-b}{\pi k}\sin kc,\qquad C_{-k}=C_{k},$
for $k=1,2,\ldots$. Now, in order to give the main results, we prove the following lemma. Without loss of generality, we assume that $\Psi_{n}(x)$ is a normalized eigenfunction corresponding to the eigenvalue $\lambda_{n}$.
@@
Proof. (a) Since $\Psi_{n}||=1$ and $\sin kx||=\sqrt{\pi}/\sqrt{2}$, by the Schwarz inequality, we have $|(q\Psi_{n},\sin kx)|\leq M\sqrt{\pi}/\sqrt{2}$. Considering the number of terms and the greatest summands of $R_{2m}(\lambda)$ in absolute value, by (46) and (47) we obtain,
@@
(b) Suppose the contrary, $(\Psi_{n},\sin nx)=0$. Since the system of root functions
$\{\sqrt{2}\sin kx/\sqrt{\pi}:k\in\mathbb{Z}^{+}\}$
of $L(0)$ forms an orthonormal basis for $L_{2}[0,\pi]$, we have the decomposition
$\dfrac{\pi}{2}\Psi_{n}=(\Psi_{n},\sin nx)\sin nx+\sum_{k\in\mathbb{Z}^{+},k\neq n}\left(\Psi_{n},\sin kx\right)\sin kx$
@@
$\sum_{k\in\mathbb{Z}^{+},k\neq n}|\left(\Psi_{n},\sin kx\right)|^{2}=\frac{\pi}{2}.$
Using the relation,
$(\lambda_{N}-n^{2})(\Psi_{N},\sin nx)=(q\Psi_{N},\sin nx)$
which is obtained from
$\Psi_{N}^{\prime\prime}(x)+q(x)\Psi_{N}(x)=\lambda_{N}\Psi_{N}(x),$
by multiplying both sides of the equality by $\sin nx$, where $\Psi_{N}(x)$ is an eigenfunction corresponding to the eigenvalue $\lambda_{N}$, the Bessel inequality and (47), we obtain
$\sum_{k\in\mathbb{Z}^{+},k\neq n}|\left(\Psi_{n},\sin kx\right)|^{2}=\sum_{k\in\mathbb{Z}^{+},k\neq n}\frac{|(q\Psi_{n},\sin kx)|^{2}}{|\lambda_{n}-k^{2}|^{2}}$
$\qquad\leq\frac{1}{(2n-1-M)^{2}}\sum_{k\in\mathbb{Z},k\neq\pm n}|(q\Psi_{n},\sin kx)|^{2}<\frac{\pi(M)^{2}}{2(2n-1-M)^{2}}<\frac{\pi}{2},$
which contradicts $\sum_{k\in\mathbb{Z}^{+},k\neq n}|\left(\Psi_{n},\sin kx\right)|^{2}=\pi/2$ and completes the proof.
Before stating the main results, we want to approximate the following series defined in Lemma 1 in a different way.
$A_{1,n}(n^{2})$ $=\sum_{\begin{subarray}{c}k=-\infty\\ k\neq 0,-2n\end{subarray}}^{\infty}\frac{C_{k}{}^{2}}{n^{2}-(n+k)^{2}}=\sum_{\begin{subarray}{c}k=-\infty\\ k\neq 0,-2n\end{subarray}}^{\infty}\frac{C_{k}^{2}}{(2n+k)(-k)}=\sum_{\begin{subarray}{c}k=-\infty\\ k\neq-2n\end{subarray}}^{-1}\frac{C_{k}^{2}}{(2n+k)(-k)}+\sum_{k=1}^{\infty}\frac{C_{k}^{2}}{(2n+k)(-k)}$
$=\sum_{\begin{subarray}{c}k=1\\ k\neq 2n\end{subarray}}^{\infty}\biggl(\frac{C_{k}^{2}}{k(2n-k)}-\frac{C_{k}^{2}}{k(2n+k)}\biggr)=\sum_{\begin{subarray}{c}k=1\\ k\neq 2n\end{subarray}}^{\infty}\frac{2C_{k}^{2}}{(2n-k)(2n+k)}.$ (48)
To this end, we introduce the integral
$I=\int_{0}^{\pi}(P(x,n)-\frac{C_{4n}}{2n}\sin 2nx)^{2}\cos 4nxdx,$ (49)
where
$P(x,n)=\int_{0}^{x}q(t)\cos 2ntdt-C_{2n}x.$
It is obvious that
$P(0,n)=P(\pi,n)=0,\qquad P^{\prime}(x,n)=q(x)\cos 2nx-C_{2n},$
and
$P(x,n)=\begin{cases}\dfrac{a}{2n}\sin 2nx-C_{2n}x&\text{if }x\in[0,c],\\[5.69054pt] \dfrac{b}{2n}\sin 2nx-C_{2n}x+C_{2n}\pi&\text{if }x\in(c,\pi].\end{cases}$
The Fourier sine coefficients of $P(x,n)$ are
$P_{k}(n)=\int_{0}^{\pi}P(x,n)\sin kxdx=\frac{\pi(C_{2n+k}+C_{2n-k})}{2k},\qquad P_{2n}(n)=\frac{\pi C_{4n}}{4n}.$
Using the Fourier decomposition
$P(x,n)-\frac{C_{4n}}{2n}\sin 2nx=\sum_{\begin{subarray}{c}k=1\\ k\neq 2n\end{subarray}}^{\infty}\frac{C_{2n+k}+C_{2n-k}}{k}\sin kx$
of $P(x,n)-\dfrac{C_{4n}}{2n}\sin 2nx$ in the integral $I$ defined by (49), we obtain
$I$ $=\int_{0}^{\pi}(P(x,n)-\frac{C_{4n}}{2n}\sin 2nx)^{2}\cos 4nxdx=\int_{0}^{\pi}\biggl(\sum_{\begin{subarray}{c}k=1\\ k\neq 2n\end{subarray}}^{\infty}\frac{C_{2n+k}+C_{2n-k}}{k}\sin kx\biggr)^{2}\cos 4nxdx$
$=\int_{0}^{\pi}\biggl(\sum_{\begin{subarray}{c}k=1\\ k\neq 2n\end{subarray}}^{\infty}\frac{C_{4n-k}+C_{k}}{2n-k}\sin(2n-k)x\frac{C_{4n+k}+C_{k}}{2n+k}\sin(2n+k)x\biggr)\cos 4nxdx$
$=-\frac{\pi}{4}\sum_{\begin{subarray}{c}k=1\\ k\neq 2n\end{subarray}}^{\infty}\frac{(C_{4n-k}+C_{k})(C_{4n+k}+C_{k})}{(2n-k)(2n+k)}$
and therefore
$\sum_{\begin{subarray}{c}k=1\\ k\neq 2n\end{subarray}}^{\infty}\frac{(C_{4n-k}+C_{k})(C_{4n+k}+C_{k})}{(2n-k)(2n+k)}=\frac{-4I}{\pi}.$ (50)
Now, the integration by parts in the integral $I$ defined by (49) gives
$I$ $=-\frac{1}{2n}\int_{0}^{\pi}P(x,n)(q(x)\cos 2nx-C_{2n})\sin 4nxdx+\frac{C_{4n}}{4n^{2}}\int_{0}^{\pi}\sin 2nx(q(x)\cos 2nx-C_{2n})\sin 4nxdx$
$=-\frac{1}{4n}\int_{0}^{\pi}P(x,n)q(x)[\sin 6nx+\sin 2nx]dx+\frac{C_{2n}}{2n}\int_{0}^{\pi}P(x,n)\sin 4nxdx+\frac{C_{4n}}{8n^{2}}\int_{0}^{\pi}q(x)[1-\cos 8nx]dx$
$=-\frac{1}{4n}I_{1}+\frac{C_{2n}}{2n}I_{2}-\frac{\pi C_{4n}C_{8n}}{8n^{2}},$
where
$I_{1}=\int_{0}^{\pi}P(x,n)q(x)[\sin 6nx+\sin 2nx]dx,\qquad I_{2}=\int_{0}^{\pi}P(x,n)\sin 4nxdx.$
By direct calculations, we obtain
$I_{1}$ $=-\frac{ab\pi}{4n}+\frac{(b^{2}-a^{2})\sin 8nc}{32n^{2}}-\frac{(b-a)^{2}}{16\pi n^{3}}+\frac{(b-a)^{2}\cos 4nc}{18\pi n^{3}}+\frac{(b-a)^{2}\cos 8nc}{144\pi n^{3}},$
$I_{2}$ $=\frac{(b-a)\sin 2nc}{16n^{2}}-\frac{(b-a)\sin 6nc}{48n^{2}},$
and
$I$ $=\frac{ab\pi}{16n^{2}}-\frac{(b^{2}-a^{2})\sin 8nc}{128n^{3}}+\frac{13(b-a)^{2}\cos 4nc}{3.2^{9}\pi n^{4}}$
$+\frac{(b-a)^{2}}{128\pi n^{4}}-\frac{5(b-a)^{2}\cos 8nc}{9.2^{7}\pi n^{4}}+\frac{(b-a)^{2}\cos 12nc}{2^{9}\pi n^{4}}.$
Therefore, by (48)-(50) we have
$|A_{1,n}(n^{2})|$ $<\frac{12}{\pi}|I|=\frac{3ab}{4n^{2}}-\frac{3(b^{2}-a^{2})\sin 8nc}{32\pi n^{3}}$ (51)
$+\frac{13(b-a)^{2}\cos 4nc}{2^{7}\pi^{2}n^{4}}+\frac{3(b-a)^{2}}{32\pi^{2}n^{4}}-\frac{5(b-a)^{2}\cos 8nc}{96\pi^{2}n^{4}}+\frac{3(b-a)^{2}\cos 12nc}{2^{7}\pi^{2}n^{4}}.$
Now, letting $m$ tend to infinity in equation (2), we obtain the following main results.
@@
If the assumptions about $M$ given in Condition 1 hold, then $\lambda_{n}$ is an eigenvalue of $L(q)$ if and only if it is the root of the equation
$\lambda-n^{2}+C_{2n}-\sum\limits_{k=1}^{\infty}a_{k}(\lambda)=0$ (52)
in the set $I_{n}:=[n^{2}-M,n^{2}+M]$.
@@
$\lambda_{n}-n^{2}+C_{2n}-\sum\limits_{k=1}^{\infty}a_{k}(\lambda_{n})=0,$
where $C_{k}=\dfrac{1}{\pi}{\displaystyle\int_{0}^{\pi}}q(x)\cos kxdx$.
Now, we prove that the root of (52) lying in the interval $I_{n}$ is an eigenvalue of the operator $L$. To do this, we first give the following relations:
@@
and
$a_{1}(n^{2})=A_{1,n}(n^{2})+B_{1,n}(n^{2}),$ (53)
where $A_{1,n}(n^{2})$ and $B_{1,n}(n^{2})$ are given by (11) and (12), respectively.
@@
$|F_{n}(\lambda)|\geq|\lambda-n^{2}|-|C_{2n}|-|A_{1,n}(n^{2})|-|B_{1,n}(n^{2})|.$
Estimating $|A_{1,n}(n^{2})|$ and $|B_{1,n}(n^{2})|$ by (51) and (44), we have
@@
for $k\geq 2$, where $a_{1}(n^{2})$ is defined by (53). Therefore, it follows by the geometric series formula that
@@
$\qquad\leq|a_{1}(\lambda)-a_{1}(n^{2})|+\sum\limits_{k=2}^{\infty}|a_{k}(\lambda)|$
$\qquad<\frac{16}{45\pi^{2}}+\frac{\pi+1}{\pi^{2}(\pi-1)}<0.1534$
@@
Now, we can use numerical methods by taking finite sums instead of the infinite series in (52) and obtain

### v2

@@ 2 On the Sharp Asymptotic Formulas
$D_{2}(n,p)=\frac{i}{4\pi n}\biggl(\int_{0}^{\pi-c}p(x)G(x,n)e^{-i2nx}\,dx+\int_{\pi-c}^{\pi}p(x)G(x,n)e^{-i2nx}\,dx\biggr).$
By direct calculations, similar to the calculations above for the terms in $D_{1}(n,q)$ and $D_{2}(n,q)$, we obtain
@@
$D_{1}(n,p)=\frac{ab}{8n^{2}}-\frac{i(b-a)(b+3a)(e^{i2nc}-1)^{2}}{32\pi n^{3}}+\frac{(b-a)^{2}(e^{i2nc}-1)^{2}}{16\pi^{2}n^{4}},$ (38)
@@
Thus, using (38), (39), Lemma 1, and simplifying a bit, we write
@@
Now, we calculate $B_{1,n}(n^{2})$ defined by (21). Expressing $Q(x)$ defined in Lemma 2 for the Kronig-Penney potential as
$Q(x)=\begin{cases}ax&\text{if }x\in[0,c]\\[5.69054pt] bx-b\pi&\text{if }x\in(c,\pi],\end{cases}$
@@
Therefore, by $C_{k}=\dfrac{a-b}{\pi k}\sin kc$, (37), (40), and (41), we can calculate:
$-C_{2n}+\frac{1}{\pi}\int_{0}^{\pi}Q^{2}(x)\cos 2nxdx+D(n,q)+2\operatorname{Re}D(n,p)$
$\qquad=-\frac{a-b}{2\pi n}\sin 2nc-\frac{ab\cos 2nc}{2n^{2}}+\frac{ab}{4n^{2}}+O\bigl(\frac{1}{n^{3}}\bigr).$
@@
$\lambda_{n}=n^{2}+\frac{b-a}{2\pi n}\sin 2nc+\frac{ab(1-2\cos 2nc)}{4n^{2}}+O\bigl(\frac{1}{n^{3}}\bigr).$
@@ 3 Estimates for the Small Eigenvalues
in the set $I_{n}:=[n^{2}-M,n^{2}+M]$, where $M=\max\{|a|,b\}$, and the infinite series $a_{k}$ is defined by (5), if the following condition holds:
@@
To consider the $n$th eigenvalue $\lambda_{n}$, for $n>1$, we assume that the condition M<(2n1)/2M<(2n-1)/2 holds. In case $n=1$, we assume that the condition $M\leq 1/2$ holds.
Then, to use numerical methods, we take finite sums instead of the infinite series in (42). To approximate the roots of equation (42), we use the fixed point iteration. Using the Banach fixed point theorem, we prove that the equation containing the finite sums has a unique solution in the appropriate set $I_{n}$ (see Theorem 4). Moreover, we give error analysis (see Theorem 4 and Theorem 5) and present a numerical example.
Now, we state some preliminary facts. It is well known that the spectrum of the operator $L(q)$ is discrete and for large enough $n$, there is one eigenvalue (counting multiplicity) in the neighborhood of $n^{2}$. See the basic and detailed classical results in [1, 2, 4, 5] and references therein. The eigenvalues of the operator $L(0)$ are $n^{2}$, for $n\in\mathbb{N}$, where $\mathbb{N}$ is the set of positive integers, and all the eigenvalues of $L(0)$ are simple.
It is also known that [6], $|\lambda_{n}-n^{2}|\leq M$, for $n\geq 1$. Therefore, we have
@@
for $n\geq 1$. If $k\neq\pm n$, then $|\lambda_{n}-k^{2}|\geq|n^{2}-k^{2}|-M=|n-k||n+k|-M\geq(2n-1)-M$, for $n\geq 1$ and under the assumptions about $M$ given in Condition 1. In particular, if $n=1$, we have $|\lambda_{1}|\leq 1+M$ and $|\lambda_{1}-k^{2}|\geq||\lambda_{1}|-k^{2}|\geq 4-|\lambda_{1}|\geq 3-M$, for $|k|\geq 2$. Besides, if $n\geq 2$, we have $|\lambda_{n}|\geq|\lambda_{2}|\geq 4-M$ and $|\lambda_{n}-k^{2}|\geq||\lambda_{2}|-k^{2}|\geq|\lambda_{2}|-1\geq 3-M$, for $k\neq\pm n$.
@@
We remind that, since the potential $q$ is of the form (9), we have $C_{k}=\dfrac{a-b}{\pi k}\sin kc$ and $C_{-k}=C_{k}$, for $k=1,2,\ldots$. Now, in order to give the main results, we prove the following lemma. Without loss of generality, we assume that $\Psi_{n}(x)$ is a normalized eigenfunction corresponding to the eigenvalue $\lambda_{n}$.
@@
Proof. (a) Since $\Psi_{n}||=1$ and $\sin kx||=\sqrt{\pi}/\sqrt{2}$, by the Schwarz inequality, we have $|(q\Psi_{n},\sin kx)|\leq M\sqrt{\pi}/\sqrt{2}$. Considering the number of terms and the greatest summands of $R_{2m}(\lambda)$ in absolute value, by (43), we obtain,
@@
(b) Suppose the contrary, $(\Psi_{n},\sin nx)=0$. Since the system of root functions $\{\sqrt{2}\sin kx/\sqrt{\pi}:k\in\mathbb{N}\}$ of $L(0)$ forms an orthonormal basis for $L_{2}[0,\pi]$, we have the decomposition
$\dfrac{\pi}{2}\Psi_{n}=(\Psi_{n},\sin nx)\sin nx+\sum_{k\in\mathbb{N},k\neq n}\left(\Psi_{n},\sin kx\right)\sin kx$
@@
$\sum_{k\in\mathbb{N},k\neq n}|\left(\Psi_{n},\sin kx\right)|^{2}=\frac{\pi}{2}$ (44)
On the other hand, using the relation $(\lambda_{N}-n^{2})(\Psi_{N},\sin nx)=(q\Psi_{N},\sin nx)$, the Bessel inequality and (43), we obtain
$\sum_{k\in\mathbb{N},k\neq n}|\left(\Psi_{n},\sin kx\right)|^{2}=\sum_{k\in\mathbb{N},k\neq n}\frac{|(q\Psi_{n},\sin kx)|^{2}}{|\lambda_{n}-k^{2}|^{2}}$
$\qquad\leq\frac{1}{(2n-1-M)^{2}}\sum_{k\in\mathbb{Z},k\neq\pm n}|(q\Psi_{n},\sin kx)|^{2}<\frac{\pi(M)^{2}}{2(2n-1-M)^{2}}<\frac{\pi}{2}.$
which contradicts Parseval’s equality (44).
Before giving the main results, we calculate the number $C_{2n}+A_{1,n}(n^{2})-B_{1,n}(n^{2})$ by (14), (37), (40), and (41):
$-C_{2n}+A_{1,n}(n^{2})-B_{1,n}(n^{2})=\frac{b-a}{2\pi n}\sin 2nc+\frac{ab(1-2\cos 2nc)}{4n^{2}}+\frac{(b-a)(\cos 2nc-1)(2a\sin 2nc+ib)}{16n^{3}}$
$\qquad+\frac{(b^{2}-a^{2})(2\sin 2nc(2+\cos 2nc)+i(\cos 2nc-1))}{16\pi n^{3}}+\frac{(b-a)^{2}\cos 2nc(\cos 2nc-1)}{16\pi^{2}n^{4}}$
$\qquad-\frac{-(b-a)[2(\sin 2nc+\sin(2n-1)c-\sin(4n-1)c)+ib(1-e^{-i2nc})(e^{i2nc}+e^{ic})]}{16n^{2}(2n-1)}$
$\qquad+\frac{(b-a)^{2}(\cos 4nc-\cos 2nc)}{32\pi n^{4}}+\frac{(b^{2}-a^{2})(e^{-i2nc}-1)-(b-a)^{2}(e^{ic}+e^{i(2n-1)c}-2\cos(4n-1)c)}{16\pi n^{2}(2n-1)^{2}}.$ (45)
Now, we state the following main result.
@@
If the assumptions about $M$ given in Condition 1 hold, then $\lambda_{n}$ is an eigenvalue of $L(q)$ if and only if it is the root of equation (42) in the set $I_{n}:=[n^{2}-M,n^{2}+M]$.
@@
$\lambda_{n}-n^{2}+C_{2n}-\sum\limits_{k=1}^{\infty}a_{k}(\lambda_{n})=0.$
Now, we prove that the root of (42) lying in the interval $I_{n}$ is an eigenvalue of the operator $L$. To do this, we first give the following relation:
@@
where $a_{1}(\lambda)$ is defined in (3), $a_{1}(n^{2})=A_{1,n}(n^{2})-B_{1,n}(n^{2})$, $A_{1,n}(n^{2})$ and $B_{1,n}(n^{2})$ are given in (11).
@@
$|F_{n}(\lambda)|\geq|\lambda-n^{2}|-|C_{2n}-A_{1,n}(n^{2})+B_{1,n}(n^{2})|.$
Estimating $|C_{2n}-A_{1,n}(n^{2})+B_{1,n}(n^{2})|$ by (45), we have
@@
for $k\geq 2$. Therefore, it follows by the geometric series formula that
@@
$\qquad\leq|a_{1}(\lambda)-a_{1}(n^{2})|+\sum\limits_{k=2}^{\infty}|a_{k}(\lambda)|<\frac{16}{45\pi^{2}}+\frac{\pi+1}{\pi^{2}(\pi-1)}<0.1534$
@@
Now, we can use numerical methods by taking finite sums instead of the infinite series in (42) and write